OpenStudy (anonymous):

Let's say we have a circle with a radius of 14 inches. What is the circumference of this circle in inches? Round your answer to the nearest whole number.

OpenStudy (anonymous):

circumference of circle is 2pi*r

OpenStudy (anonymous):

C = 2*PI*R --> R is the radius, and Pi is the constant approximately 3.14. Plug in the given values in to the equation and solve. C = 2*Pi*(14) = 28*Pi = 88 inches (approximately, rounded off to the nearest whole number)

OpenStudy (zpupster):

if your dealing with circles you should memorize some formulas Circumference = 2pi r = diameter(pi) Area = (pi) r^2 these will come in handy for circle problems